Solve : d2y/dx2+a2y=sec ax
1
Expert's answer
2020-11-19T18:25:58-0500

"\displaystyle \frac{\mathrm{d}^2y}{\mathrm{d}x^2} + a^2y = \sec(ax)

\\ \textsf{The solution to the above equation is}
\\ y = y_c + y_p \\
\textsf{where}\, y_c \, \textsf{is the complementary factor and}
\\ y_p\, \textsf{ is the particular integral}
\\ \textsf{The auxiliary equation is} \,m^2 + a^2 = 0 \\
 
\begin{aligned}
 (m - ja)(m + ja) &= 0 \\ 
\therefore m &= \pm ja \hspace{1cm} \{\textsf{where}\, j\, \textsf{is a complex number}\}
\end{aligned} \\

\textsf{Recall that if the solution of the auxiliary equation}\\
\textsf{of a second-order differential equation is}\\
m = \alpha \pm j\beta, \textsf{the general solution is} \\

y = e^{\alpha x}(C_1\cos{\beta x} + C_2\sin{\beta x})\\

\therefore y_c = C_1\cos(ax) + C_2\sin(ax)\\

\textsf{The Wronskian of the two solutions is} \\
W(x) = \begin{vmatrix} 
\cos(ax) & \sin(ax)\\ 
\frac{\mathrm{d}}{\mathrm{d}x}(\cos(ax)) & \frac{\mathrm{d}}{\mathrm{d}x}(\sin(ax)) 
\end{vmatrix} = 

\begin{vmatrix} 
\cos(ax) & \sin(ax)\\
 -a\sin(ax) & a\cos(ax)
 \end{vmatrix} \\ 

\begin{aligned}
&= a\cos^2(ax) + a\sin^2(ax)\\
&= a(\cos^2(ax) + \sin^2(ax)) \\
&= a
\end{aligned}\\

\therefore\textsf{ Our particular solution will be given by}\\ 
y_p = V_1(x)\cos(ax) + V_2(x)\sin(ax)\\

\textsf{Where}\, V_1(x) = -\int \frac{r(x)\sin(ax)}{W(x)} \, \mathrm{d}x,\, V_2(x) = \int \frac{r(x)\cos(ax)}{W(x)}\, \mathrm{d}x \, \textsf{and} \, \\

\begin{aligned} 
V_1(x) &= -\int \frac{r(x)\sin(ax)}{W(x)} \, \mathrm{d}x \\
&= -\int \frac{\sec(ax)\sin(ax)}{a}\, \mathrm{d}x\\
&= -\int \frac{\tan(ax)}{a}\, \mathrm{d}x\\
&= -\frac{\ln(\sec(ax))}{a^2} + C
\end{aligned} \\

\begin{aligned} 
V_2(x) &= \int \frac{r(x)\cos(ax)}{W(x)}\, \mathrm{d}x \\
&= \int \frac{\sec(ax)\cos(ax)}{a}\, \mathrm{d}x\\
&= \int \frac{1}{a}\, \mathrm{d}x\\
&= \frac{x}{a} + C
\end{aligned} \\
\textsf{The constant terms of the integration can be}
\\\textsf{ignored since we are trying to find a non-constant}
\\\textsf{solution to the differential equation}

\\ \begin{aligned}
\therefore y &= y_c + y_p \\
&=C_1\cos(ax) + C_2\sin(ax) - \frac{1}{a^2}\ln(\sec(ax))\cos(ax) +\frac{x}{a}\sin(ax)
\end{aligned}"